# Determination of the stably free cancellation property for orders

### v1

@@ 5.5. A criterion assuming the Eichler condition for one component
Let $I$ and $J$ be two-sided ideals of $\Lambda$ such that $I\cap J=0$ and $K(I+J)=A$. Let $\Lambda_{1}=\Lambda/I$, let $\Lambda_{2}=\Lambda/J$ and let $\overline{\Lambda}=\Lambda/(I+J)$. Assume that $A_{1}:=K\Lambda_{1}\cong KJ$ satisfies the Eichler condition. For $i=1,2$, let $g_{i}:\Lambda_{i}\rightarrow\overline{\Lambda}$ be the canonical surjection. Let $\partial$ and $W$ be as in Corollary 5.3. Then $\overline{\Lambda}$ is finite and we have
@@
Assume the setting of Proposition 5.11. Then $W$ is a group. Suppose that $N$ is a subgroup of $W$ containing $[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. Then $N\trianglelefteq\ker(\partial)$ and $\ker(\partial)/N$ is a finite abelian group. Let $S$ be a generating subset of $\ker(\partial)/N$. Suppose that $\Lambda_{2}$ has SFC. Then the following assertions are equivalent.
@@ 6.1. A fibre product for integral group rings
We first record the following well-known instance of a fibre product. Let $G$ be a finite group, let $N$ be a normal subgroup and let $H=G/N$. Let $n=|N|$ and let $\mathrm{Tr}_{N}=\sum_{n\in N}n\in\mathbb{Z}[G]$ be the trace element associated to $N$. Let $\pi_{1}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[H]$ be the canonical projection and let $I=\ker(\pi_{1})$. Let $J=\mathrm{Tr}_{N}\mathbb{Z}[G]$, let $\Gamma=\mathbb{Z}[G]/J$ and let $\pi_{2}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[G]/J$ be the canonical projection. Then $I\cap J=0$ and $I+J=n\mathbb{Z}[G]$. Thus, by Example 5.1 we have a fibre product
@@ 6.2. Integral group rings of the form $\mathbb{Z}[H\times C_{2}]$
Let $\tilde{T}$ denote the binary tetrahedral group. Then $\tilde{T}\cong\mathrm{SL}_{2}(\mathbb{F}_{3})\cong G_{(24,3)}$.
@@
The integral group ring $\mathbb{Z}[\tilde{T}\times C_{2}]$ has SFC.
@@
The algorithm of [BB06, §3] shows that $|\mathrm{Cl}(\mathbb{Z}[\tilde{T}])|=2$ and $|\mathrm{Cl}(\mathbb{Z}[\tilde{T}\times C_{2}])|=2^{6}$. By Remark 5.9, we have $|\mathbb{F}_{2}[\tilde{T}]^{\times}|=|\mathbb{F}_{2}[\tilde{T}]||(\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}]))^{\times}|$, where $J(\mathbb{F}_{2}[\tilde{T}])$ denotes the Jacobson radical of $\mathbb{F}_{2}[\tilde{T}]$. By [MS22, Theorem 2.5] we have $\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}])\cong\mathbb{F}_{2}\oplus\mathbb{F}_{4}$ (this can also be verified using the algorithm of [FR85]). Hence $|J(\mathbb{F}_{2}[\tilde{T}])|=2^{24-3}=2^{21}$. Moreover, $(\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}]))^{\times}\cong(\mathbb{F}_{2}\oplus\mathbb{F}_{4})^{\times}\cong C_{3}$. Therefore $|\mathbb{F}_{2}[\tilde{T}]^{\times}|=3\cdot 2^{21}$.
@@
To prove equality, we consider the following units of integral groups rings as introduced by Ritter and Segal in [RS91]. For $g,h\in\tilde{T}$, let $d=\operatorname{ord}(g)$ and define
@@
and note that this is a subgroup of $V$. By Remark 6.3, we obtain an explicit embedding of $\mathbb{F}_{2}[\tilde{T}]^{\times}$ into $\mathrm{GL}_{24}(\mathbb{F}_{2})$. Since a calculation on a computer algebra system such as Magma [BCP97] or Oscar [OSC24, DEF+24] shows that the image of $U$ under this embedding is a subgroup of $\mathrm{GL}_{24}(\mathbb{F}_{2})$ of order $3\cdot 2^{17}$, we must have $U=V$ and $|V|=3\cdot 2^{17}$. Therefore we conclude that $\mathbb{Z}[\tilde{T}\times C_{2}]$ has SFC. ∎
@@
Using Theorem 6.6 as an input, Nicholson [Nic24, Theorem 1.7] has shown that the integral group ring $\mathbb{Z}[\tilde{T}\times C_{2}]$ in fact has LFC.
@@
By [Oli81, Theorem 3], for any finite group $G$ and any prime $p$, the Sylow $p$-subgroup of $SK_{1}(\mathbb{Z}[G])$ is generated by induction from the $p$-elementary subgroups of $G$. If $H$ is any abelian subgroup of $\tilde{T}\times C_{2}$ then $H\cong C_{2}\times C_{2}$, $C_{2}\times C_{4}$, $C_{2}\times C_{6}$ or $H$ is cyclic and so $SK_{1}(\mathbb{Z}[H])=0$ by [ADOS87, Theorem 4.9]. If $H$ is any non-abelian elementary subgroup of $\tilde{T}\times C_{2}$ then $H\cong Q_{8}$ or $Q_{8}\times C_{2}$ and so $SK_{1}(\mathbb{Z}[H])=0$ by [Oli81, Theorem 4] and [Oli88, Example 9.9(ii)], respectively. Therefore $SK_{1}(\mathbb{Z}[\tilde{T}\times C_{2}])=0$. ∎
We now have the following consequence of Theorem 6.6, which has also been shown independently by Nicholson [Nic24, Proposition 4.14] .
@@
Let $H=\tilde{T}\times C_{2}$. Then $SK_{1}(\mathbb{Z}[H])=0$ by Lemma 6.8 and it is straightforward to deduce from the character table of $H$ that $r_{\mathbb{R}}(H)=r_{\mathbb{Q}}(H)=10$. Thus $K_{1}(\mathbb{Z}[H])$ is represented by (trivial) units by Proposition 4.12. We also have that $m_{\mathbb{H}}(H)=2$, which can be deduced from the character table of $H$ along with its Frobenius-Schur indicators. Thus the desired result now follows from Theorems 4.9 and 6.6. ∎
@@ 6.3. A useful reduction criterion for certain integral group rings
The following reduction criterion only uses the sizes of certain locally free class groups, and thus can be applied relatively easily.
@@ 7.2. A criterion for freeness
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Let $u\in\mathcal{M}^{\times}$ and let $\alpha=u\beta$. Then $\overline{\alpha},\overline{\beta},\overline{u}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and the following are equivalent:
@@
[BHJ22, Proposition 4.3 and Theorem 4.5] only assumed that $\beta\in\mathcal{M}$ but should have assumed that $\beta\in\mathcal{M}\cap A^{\times}$, or equivalently, that $\beta\in\mathcal{M}$ and that $X$ is a full left ideal of $\Lambda$, that is, $KX=A$. Note that any locally free left ideal of $\Lambda$ is necessarily full. Moreover, if $\Lambda\cap V\neq\mathfrak{f}\cap V$ for all non-trivial $K$-vector subspaces $V$ of $A$, then the condition $X+\mathfrak{f}=\Lambda$ also implies that $X$ is full. The strengthened assumption is needed to ensure that the condition $X=\Lambda\alpha$ implies that $X$ is in fact free over $\Lambda$ (see [BHJ22, Lemma 2.3]); without this, the implication (i) $\Rightarrow$ (v) in Proposition 7.1 may not hold. This omission does not affect any of the other results of [BHJ22].
@@
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then the following are equivalent:
@@
    $X$ is free over $\Lambda$,
@@
    there exists $u\in\mathcal{M}^{\times}$ and $\overline{\alpha}\in(\Lambda/\mathfrak{f})^{\times}$ such that $\bar{u}=\bar{\beta}\bar{\alpha}$,
@@
    there exists $u\in\mathcal{M}^{\times}$ such that $\pi(\bar{\beta})=\pi(\bar{u})$.
@@
This follows from Proposition 7.1 after substituting $u$ and $\bar{\alpha}$ for their inverses. ∎
@@ 7.4. Decomposition via central idempotents
is a decomposition of $A$ into indecomposable two-sided ideals (see [CR81, (3.22)]). Each $A_{i}$ is a simple $K$-algebra with identity element $e_{i}$. The centres $K_{i}:=Z(A_{i})$ are finite field extensions of $K$ via $K\rightarrow K_{i}$, $\alpha\mapsto\alpha e_{i}$, and we have $K$-algebra isomorphisms $A_{i}\cong\mathrm{Mat}_{n_{i}}(D_{i})$ where $D_{i}$ is a skew field with $Z(D_{i})\cong K_{i}$ (see [CR81, (3.28)]). The Wedderburn decomposition (7.4) induces decompositions
@@ 7.6. Idèles
$J(A)=\{(a_{\mathfrak{p}})_{\mathfrak{p}}\in\textstyle{\prod_{\mathfrak{p}}}A^{\times}_{\mathfrak{p}}\mid a_{\mathfrak{p}}\in\widehat{\Lambda}^{\times}_{\mathfrak{p}}\text{ for almost all }\mathfrak{p}\}.$
@@
involving various completions, where, for given $i\in\{1,\ldots,r\}$, $\mathfrak{P}$ runs through all maximal ideals of $\mathcal{O}_{K_{i}}$ dividing $\mathfrak{p}$ and $A_{i,\mathfrak{P}}$ is defined as $(A_{i})_{\mathfrak{P}}$. Using the above isomorphism, we will often write elements of $J(A)$, resp. $A_{\mathfrak{p}}$, as tuples $(a_{i,\mathfrak{P}})_{i,\mathfrak{P}}$, where $\mathfrak{P}$ ranges over all maximal ideals of $\mathcal{O}_{K_{i}}$, resp. over those that contain $\mathfrak{p}$. Similarly, we denote by $J(C)$ the group of idèles of $C$. Again one has a canonical isomorphism
@@
7.7. Review of results of Bley-Boltje
@@
Let $x\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ and let $a\in\Lambda\cap A^{\times}$ such that $a\equiv x\pmod{\mathfrak{f}}$. Then $\nu(x)=(\mathrm{nr}(a)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})$ in $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$.
Let $x\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$ and let $a\in\mathcal{M}\cap A^{\times}$ such that $a\equiv x\pmod{\mathfrak{f}}$. Then define
$\tilde{\nu}(x)=(\mathrm{nr}(a)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})\in I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}.$ (7.9)

### v2

@@ 5.5. A criterion assuming the Eichler condition for one component
Let $I$ and $J$ be two-sided ideals of $\Lambda$ such that $K(I+J)=A$ and $I\cap J=0$. Let $\Lambda_{1}=\Lambda/I$, let $\Lambda_{2}=\Lambda/J$ and let $\overline{\Lambda}=\Lambda/(I+J)$. Assume that $A_{1}:=K\Lambda_{1}\cong KJ$ satisfies the Eichler condition. For $i=1,2$, let $g_{i}:\Lambda_{i}\rightarrow\overline{\Lambda}$ be the canonical surjection. Let $\partial$ and $W$ be as in Corollary 5.3. Then $\overline{\Lambda}$ is finite and we have
@@
Assume the setting of Proposition 5.11. Then $W$ is a group. Suppose that $N$ is a subgroup of $W$ containing $[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. Then $N\trianglelefteq\ker(\partial)$ and $\ker(\partial)/N$ is a finite abelian group. Let $S$ be a generating subset of $\ker(\partial)/N$. Suppose that $\Lambda_{2}$ has SFC. Then the following assertions are equivalent:
@@ 6.1. A fibre product for integral group rings
We first record the following well-known instance of a fibre product. Let $G$ be a finite group, let $N$ be a normal subgroup and let $H=G/N$. Let $n=|N|$ and let $\mathrm{Tr}_{N}=\sum_{\tau\in N}\tau\in\mathbb{Z}[G]$ be the trace element associated to $N$. Let $\pi_{1}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[H]$ be the canonical projection and let $I=\ker(\pi_{1})$. Let $J=\mathrm{Tr}_{N}\mathbb{Z}[G]$, let $\Gamma=\mathbb{Z}[G]/J$ and let $\pi_{2}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[G]/J$ be the canonical projection. Then $I\cap J=0$ and $I+J=n\mathbb{Z}[G]$. Thus, by Example 5.1 we have a fibre product
@@ 6.2. Integral group rings of the form $\mathbb{Z}[H\times C_{2}]$
Let $\tilde{T}\cong\mathrm{SL}_{2}(\mathbb{F}_{3})\cong G_{(24,3)}$ denote the binary tetrahedral group. Then $\tilde{T}\times C_{2}\cong G_{(48,32)}$.
@@
The group ring $\mathbb{Z}[\tilde{T}\times C_{2}]$ has SFC.
@@
The algorithm of [BB06, §3] shows that $|\mathrm{Cl}(\mathbb{Z}[\tilde{T}])|=2$ and $|\mathrm{Cl}(\mathbb{Z}[\tilde{T}\times C_{2}])|=2^{6}$. By Remark 5.9, we have $|\mathbb{F}_{2}[\tilde{T}]^{\times}|=|J(\mathbb{F}_{2}[\tilde{T}])||(\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}]))^{\times}|$, where $J(\mathbb{F}_{2}[\tilde{T}])$ denotes the Jacobson radical of $\mathbb{F}_{2}[\tilde{T}]$. By [MS22, Theorem 2.5] we have $\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}])\cong\mathbb{F}_{2}\oplus\mathbb{F}_{4}$ (this can also be verified using the algorithm of [FR85]). Hence $|J(\mathbb{F}_{2}[\tilde{T}])|=2^{24-3}=2^{21}$. Moreover, $(\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}]))^{\times}\cong(\mathbb{F}_{2}\oplus\mathbb{F}_{4})^{\times}\cong C_{3}$. Therefore $|\mathbb{F}_{2}[\tilde{T}]^{\times}|=3\cdot 2^{21}$.
@@
To prove equality, we consider the following units of integral group rings as introduced by Ritter–Segal [RS91]. For $g,h\in\tilde{T}$, let $d=\operatorname{ord}(g)$ and define
@@
and note that this is a subgroup of $V$. By Remark 6.3, we obtain an explicit embedding of $\mathbb{F}_{2}[\tilde{T}]^{\times}$ into $\mathrm{GL}_{24}(\mathbb{F}_{2})$. Since a calculation on a computer algebra system such as Magma [BCP97] or Oscar [OSC24, DEF+24, FHHJ17] shows that the image of $U$ under this embedding is a subgroup of $\mathrm{GL}_{24}(\mathbb{F}_{2})$ of order $3\cdot 2^{17}$, we must have $U=V$ and $|V|=3\cdot 2^{17}$. Therefore we conclude that $\mathbb{Z}[\tilde{T}\times C_{2}]$ has SFC. ∎
@@
Using Theorem 6.6 as an input, Nicholson [Nic24b, Theorem 1.7] has shown that the integral group ring $\mathbb{Z}[\tilde{T}\times C_{2}]$ in fact has LFC.
@@
By [Oli81, Theorem 3], for any finite group $G$ and any prime $p$, the Sylow $p$-subgroup of $SK_{1}(\mathbb{Z}[G])$ is generated by induction from the $p$-elementary subgroups of $G$. If $H$ is any abelian subgroup of $\tilde{T}\times C_{2}$ then $H\cong C_{n}$ or $C_{2}\times C_{n}$ for some $n$, and so $SK_{1}(\mathbb{Z}[H])=0$ by [ADOS87, Theorem 4.9]. If $H$ is any non-abelian elementary subgroup of $\tilde{T}\times C_{2}$ then $H\cong Q_{8}$ or $Q_{8}\times C_{2}$ and so $SK_{1}(\mathbb{Z}[H])=0$ by [Oli81, Theorem 4] and [Oli88, Example 9.9(ii)], respectively. Therefore the desired result holds. ∎
We now have the following consequence of Theorem 6.6, which has also been shown independently by Nicholson [Nic24b, Proposition 4.14].
@@
Let $H=\tilde{T}\times C_{2}$. Then $SK_{1}(\mathbb{Z}[H])=0$ by Lemma 6.8 and it is straightforward to deduce from the character table of $H$ that $r_{\mathbb{R}}(H)=r_{\mathbb{Q}}(H)=10$. Thus $K_{1}(\mathbb{Z}[H])$ is represented by (trivial) units by Proposition 4.14. We also have that $m_{\mathbb{H}}(H)=2$, which can be deduced from the character table of $H$ along with its Frobenius–Schur indicators. Thus the desired result now follows from Theorems 4.11 and 6.6. ∎
@@ 6.3. A useful reduction criterion for certain integral group rings
The following reduction criterion only uses the sizes of certain locally free class groups, and thus can be applied relatively easily. It will be used in the proof of Theorem 13.1 (i.e. Theorem 1.3).
@@ 7.2. A criterion for freeness
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Let $u\in\mathcal{M}^{\times}$ and let $\alpha=u\beta$. Then $\overline{\alpha},\overline{\beta},\overline{u}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and the following assertions are equivalent:
@@
[BHJ22, Proposition 4.3 and Theorem 4.5] only assumed that $\beta\in\mathcal{M}$ but should have assumed that $\beta\in\mathcal{M}\cap A^{\times}$, or equivalently, that $\beta\in\mathcal{M}$ and that $X$ is a full left ideal of $\Lambda$, that is, $KX=A$. Note that any locally free left ideal of $\Lambda$ is necessarily full. The strengthened assumption is needed to ensure that the condition $X=\Lambda\alpha$ implies that $X$ is in fact free over $\Lambda$ (see [BHJ22, Lemma 2.3]); without this, the implication (i) $\Rightarrow$ (v) in Proposition 7.1 may not hold. This omission does not affect any of the other results of [BHJ22].
@@
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then the following assertions are equivalent:
@@
    $X$ is free over $\Lambda$.
@@
    There exists $u\in\mathcal{M}^{\times}$ and $\overline{\alpha}\in(\Lambda/\mathfrak{f})^{\times}$ such that $\overline{u}=\overline{\beta}\overline{\alpha}$.
@@
    There exists $u\in\mathcal{M}^{\times}$ such that $\pi(\overline{\beta})=\pi(\overline{u})$.
@@
This follows from Proposition 7.1 after substituting $u$ and $\overline{\alpha}$ for their inverses. ∎
@@ 7.4. Decomposition via central idempotents
is a decomposition of $A$ into indecomposable two-sided ideals (see [CR81, (3.22)]). Each $A_{i}$ is a simple $K$-algebra with identity element $e_{i}$. The centres $K_{i}:=Z(A_{i})$ are finite field extensions of $K$ via $K\rightarrow K_{i}$, $\alpha\mapsto\alpha e_{i}$, and we have $K$-algebra isomorphisms $A_{i}\cong\mathrm{Mat}_{n_{i}}(D_{i})$ where $D_{i}$ is a division algebra with $Z(D_{i})\cong K_{i}$ (see [CR81, (3.28)]). The Wedderburn decomposition (7.4) induces decompositions
@@ 7.6. Idèles
$J(A)=\{(a_{\mathfrak{p}})_{\mathfrak{p}}\in\textstyle{\prod_{\mathfrak{p}}}A^{\times}_{\mathfrak{p}}\mid a_{\mathfrak{p}}\in\widehat{\Lambda}^{\times}_{\mathfrak{p}}\text{ for all but finitely many }\mathfrak{p}\}.$
@@
involving various completions, where, for given $i\in\{1,\ldots,r\}$, $\mathfrak{P}$ runs through all maximal ideals of $\mathcal{O}_{K_{i}}$ dividing $\mathfrak{p}$ and $A_{i,\mathfrak{P}}$ is defined as $(A_{i})_{\mathfrak{P}}$. Using the above isomorphism, we will often write elements of $J(A)$, resp. $A_{\mathfrak{p}}$, as tuples $(a_{i,\mathfrak{P}})_{i,\mathfrak{P}}$, where $\mathfrak{P}$ ranges over all maximal ideals of $\mathcal{O}_{K_{i}}$, resp. over those that contain $\mathfrak{p}$. Similarly, we denote by $J(C)$ the group of idèles of $C$. Again, we have a canonical isomorphism
@@
7.7. Review of results of Bley–Boltje
@@
Let $x\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ and let $a\in\Lambda\cap A^{\times}$ such that $x\equiv a\bmod{\mathfrak{f}}$. Then $\nu(x)=(\mathrm{nr}(a)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})$ in $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$.
Let $x\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$ and let $a\in\mathcal{M}\cap A^{\times}$ such that $x\equiv a\bmod{\mathfrak{f}}$. Then define
$\tilde{\nu}(x):=(\mathrm{nr}(a)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})\in I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}.$ (7.9)